OpenStudy (anonymous):

write an eqaution with a variable on both sides of the equal sign that has a infinite number of solutions. solve the equation and explain why it has in infinite number of solutions.

OpenStudy (anonymous):

make sure that the expression on one side is identical to the expression on the other

OpenStudy (anonymous):

the simplest example (which your teacher will not like) would be \(x=x\) but you can modify it a bit to obscure the fact that they are the same

OpenStudy (anonymous):

for example, since \(2(x-5)=2x-10\) by the distributive property, you could write \[2(x-5)=2x-10\] or add \(7\) so both sides and write \[2(x-5)+7=2x-3\] or whatever else you can think of